### v3

# BotArtist: Generic approach for bot detection in Twitter via semi-automatic machine learning pipeline

Abstract
Twitter, as one of the most popular social networks, provides a platform for communication and online discourse. Unfortunately, it has also become a target for bots and fake accounts, resulting in the spread of false information and manipulation. In this paper, we introduce a semi-automatic machine learning pipeline (SAMLP) designed to address the challenge of detecting Twitter bots. Through this pipeline, we have developed a comprehensive bot detection model named BotArtist, based on user profile features. SAMLP leverages nine distinct publicly available datasets to train the BotArtist model. To assess BotArtist’s performance against current state-of-the-art solutions, we have selected 34 existing bot detection methods, each utilizing a diverse range of features. Our comparative evaluation of BotArtist and these existing methods, conducted across nine public datasets under standardized conditions, reveals that the proposed model outperforms existing solutions by almost 10% in terms of F1-score achieving an average score of 83.19 and 68.5 over specific and general approaches respectively. Following the acceptance of this paper, we will make both the developed SAMLP pipeline and the BotArtist model publicly accessible.
Introduction
Online social media has become an essential part of everyday life. During the past decade, online social platforms have managed to transform the communication routine of our daily lives. Due to their growing popularity, online social media gained millions of daily active users that not only consume the information but also create a space for content creators. The main reason behind online social media is real-time access to unlimited information, where registered users can share their comments and personal opinions about popular topics. Such high interest in online social phenomena generates the opportunity and the need for different categories of human interactions and content-sharing platforms. These opportunities lead to the creation of different online social platforms each of which provides a unique user experience, with a very similar goal of real-time human communication and content sharing. Twitter, one of the most popular social networks, with millions of active users, is used for news dissemination, political discussions, and social interactions. However, the platform has also been plagued by the presence of bots and fake accounts used to manipulate and spread false information. According to the research community, the usage of manipulation techniques implemented with the use of bot accounts is registered during diverse popular topic discussions. More specifically, studies show that bot accounts are involved in discussions around the 2016 and 2020 US Presidential elections (Golovchenko et al. 2020; Badawy, Ferrara, and Lerman 2018; Howard, Kollanyi, and Woolley 2016; Shevtsov et al. 2022, 2023b). Besides the US elections a high bot activity with spreading of misleading information is also detected during election periods (presidential/parliamentary/state) in countries like Germany, Sweden, France, Spain, etc. (Neudert, Kollanyi, and Howard 2017; Pastor-Galindo et al. 2020; Bradshaw et al. 2017; Fernquist, Kaati, and Schroeder 2018; Castillo et al. 2019; Rossi et al. 2020). Furthermore, election discourse is not the only point of interest for bot account creators. Spreading propaganda, advertisement, and fake news are identified during the vaccination debate (Broniatowski et al. 2018), the advertisement of e-cigarette (Allem and Ferrara 2016; Allem et al. 2017) and more recent examples of COVID-19 pandemic (Shahi, Dirkson, and Majchrzak 2021; Ferrara 2020; Yang et al. 2021). This high activity of bot accounts raises concerns in the research community and online social media platforms about the integrity of the shared information. Twitter has been taking measures to detect and suspend these bots to maintain the integrity of the platform. Most of the published studies utilize advanced ML techniques to analyze the content and features (structural/network).
Currently, the research community offers a variety of machine learning and neural network-based bot detection methods. Each of these provided methods excels at addressing specific bot detection scenarios, yielding optimal solutions for particular use cases. Unfortunately, the performance of these provided methods significantly diminishes in more general bot detection scenarios, encompassing different periods, discussion topics, and languages, among others. Recent studies have demonstrated that existing methods still fall short of achieving flawless Twitter bot detection, regardless of the number of user characteristics they incorporate.
In our research, we address this challenge by developing a semi-automatic machine learning pipeline (SAMLP) for constructing a generic bot detection model. Next, we compare this developed model with 34 currently available state-of-the-art approaches using nine publicly available datasets. This research yields two distinct evaluation scenarios: one specific to each dataset, where each model is trained and tested on each dataset separately, and a more generic approach where the models are trained and tested on the combined dataset comprising all nine unique datasets.
As a result of our research, we introduce BotArtist, a finely tuned general bot detection model that achieves the highest F1-score performance on three out of the nine datasets, with an average F1-score performance of 83.19. Additionally, the presented approach achieves greater accuracy in the general evaluation, with an overall F1-score improvement of more than 9% and an average improvement of 5% compared to the best-performing existing bot detection approaches. Moreover, BotArtist relies on only a limited set of profile features, allowing for the prediction of historical data independently of the Twitter API.
Finally, we offer insights into the decision-making process of BotArtist through the use of the SHAP method. This approach enables us to explain the model’s decisions and highlight the distinctions between normal and bot accounts as identified by the developed model. Upon paper acceptance, we will make the SAMLP pipeline and the generated BotArtist model publicly accessible.
Related Work
Bot detection on Twitter poses a formidable challenging task, primarily due to the increasing sophistication of these bots. Research efforts in this domain can generally be categorized into three main approaches: feature-based, text-based, and graph-based techniques. Each category offers a distinct angle on bot detection, leveraging various extracted characteristics of users and their activities on the social media platform.
Feature based
Methods falling into this category perform feature engineering based on information extracted from Twitter user profiles and their activity patterns. Researchers utilize traditional machine learning or neural network classification algorithms for bot detection. These approaches employ different sets of user characteristics as feature sets, including user metadata (Kudugunta and Ferrara 2018), tweets (Miller et al. 2014), description (Hayawi et al. 2022), temporal patterns (Mazza et al. 2019), and follow relationships (Feng et al. 2021a). Moreover, some aim to enhance the scalability of feature-based methods (Yang et al. 2020), discover unknown bot accounts using correlation-based techniques (Chavoshi, Hamooni, and Mueen 2016), and improve the trade-off between precision and recall in bot detection (Morstatter et al. 2016). However, creators of bot accounts are increasingly aware of the features utilized by the research community, leading to novel bot implementations designed to evade detection based on known features (Cresci 2020). Consequently, existing feature-based methods face several challenges in the accurate detection of these new bot accounts (Feng et al. 2021a).
Text based
Text-based methods rely on natural language processing (NLP) techniques for Twitter bot detection, extracting characteristics from posted content (tweets) and user profile descriptions. Approaches in this category include word embeddings (Wei and Nguyen 2019), recurrent neural networks (RNN) (Kudugunta and Ferrara 2018), attention mechanisms (Feng et al. 2021a), and the adoption of pre-trained language models for encoding tweets (Dukić, Keča, and Stipić 2020). Researchers have also combined tweet representations with user profile characteristics (Cai, Li, and Zengi 2017), employed unsupervised machine learning models (Feng et al. 2021a), and addressed multilingual context issues (Knauth 2019). Despite the extensive existing research studies and impressive performances in text-based approaches, novel bot accounts can still evade detection by sharing stolen content from genuine users (Cresci 2020). Additionally, recent work has demonstrated that relying solely on textual information is insufficient for robust and accurate bot detection (Feng et al. 2021c).
Graph based
The graph-based category of bot detection methods combines geometric deep neural networks with graph analytics. Current implementations leverage techniques such as node centrality (Dehghan et al. 2023), node representation learning (Pham et al. 2022), graph neural networks (GNNs) (Ali Alhosseini et al. 2019), and heterogeneous GNNs (Feng et al. 2021c) to conduct graph-based bot detection. In recent research, authors have borrowed ideas from other categories to merge different approaches, such as combining graph and text-based methods (Guo et al. 2021b; Yang, Ferrara, and Menczer 2022). Furthermore, novel GNN architectures have been proposed to exploit the heterogeneity in the Twitter network (Feng et al. 2022a). Generally, these approaches hold significant promise for Twitter bot detection.
Unfortunately, existing Twitter bot detection methods have their limitations. Simplified feature-based approaches struggle to generalize and provide robust results, while more complex methods based on text or graph characteristics require intensive computational resources and large datasets for model development. These limitations, coupled with the recent announcement of the monetizing Twitter API (Twitter 2023), consist of many existing methods expensive to maintain or nearly impossible to operate daily. In this research, we address these challenges and present robust and accurate Twitter bot detection methods that rely on a lightweight set of features. Our methods require only a single user profile object request and can predict users using historical user objects (Twitter API v1.1 or v2) without additional Twitter API requests, thereby significantly reducing operational costs.
Datasets
For this research paper, we have collected nine well-known publicly available datasets (Feng et al. 2021b). All the selected datasets already contain ground truth labels, primarily obtained through manual analysis or crowd-sourcing. To simplify, we have labeled the selected datasets as follows: C-15 (Cresci et al. 2015), G-17 (Gilani et al. 2017), C-17 (Cresci et al. 2017b, a), M-18 (Yang et al. 2020), C-S-18 (Cresci et al. 2018, 2019), C-R-19 (Mazza et al. 2019), B-F-19 (Yang et al. 2019), TwiBot-20 (Feng et al. 2021b), and TwiBot-22 (Feng et al. 2022b). In table: 1, we present the information provided in each dataset, along with the volume of normal and bot accounts.
Dataset C-15 G-17 C-17 M-18 C-S-18 C-R-19 B-F-19 TwiBot-20 TwiBot-22
# Total User 5,301 2,484 14,368 50,538 13,276 693 518 229,580 1,000,000
# Human 1,950 1,394 3,474 8,092 6,174 340 380 5,237 860,057
# Bot 3,351 1,090 10,894 42,446 7,102 353 138 6,589 139,943
# Total Tweet 2,827,757 0 6,637,615 0 0 0 0 33,488,192 88,217,457
# Human Tweet 2,631,730 0 2,839,361 0 0 0 0 33,488,192 81,250,102
# Bot Tweet 196,027 0 3,798,254 0 0 0 0 33,488,192 6,967,355
# Graph Edges 7,086,134 0 6,637,615 0 0 0 0 33,488,192 170,185,937
Table 1: Description of the selected datasets and the information contained in those datasets.
Figure 1: Machine learning pipeline used for BotArtist model creation. Includes feature selection and model fine-tuning; each step is executed during separate K-Fold cross-validation.
Methodology
The proposed approach is based on the development of a semi-automatic machine learning pipeline (SAMLP). We have chosen this implementation due to its simplicity for further usage and its ability to avoid many trivial mistakes during data processing, feature selection, hyper-parameter fine-tuning, and model evaluation. The steps of the developed pipeline are presented in figure: 1, where the initial step involves data splitting.
During data splitting, it is crucial to maintain a class balance between the training/validation and testing data portions. For this purpose, we utilize a stratified data split with a ratio of 70% for training/validation and 30% for the testing data portions, respectively. This data separation allows us to keep exactly 70% of each class during the training, while preserving the natural class distribution differences, present in the dataset. The testing data portion is kept hidden and is only utilized during the final testing of the model.
Feature Selection
After the data splitting, our pipeline applies a feature selection procedure over the train/validation data portion. For this purpose, we use the K-Fold cross-validation approach, which allows us to utilize the same data portion for multiple trainings and evaluations. Although K-Fold cross-validation may lead to slightly over-optimistic performance estimations, in our case, it is not relevant since we use this procedure for alpha hyper-parameter fine-tuning of the Lasso model, and we are interested in identifying the best alpha parameter.
Additionally, since the Lasso regression model does not perform well over imbalanced datasets, we also take into account the class imbalance of the train/validation dataset. When the class balance is not equal, we utilize under-sampling of the majority class, making the data perfectly balanced and improving the prediction performance of the Lasso model. Due to the high-class imbalance and the under-sampling of the majority class, there is a very high probability of losing some important samples, which may lead to inaccurate feature selection. To address this, we utilize 10 executive repetitions to capture as many samples of the majority class as possible.
During these repetition rounds, we store the best alpha parameters based on the square mean error metric from the entire stratified k-Fold cross-validation results, where $K=5$. At the end of the procedure, we compute the most frequent best $\alpha$ parameter and train the Lasso model with the entire train/validation dataset, keeping features with non-zero coefficients. Additionally, we check if the selected $\alpha$ parameter falls within the defined limits of the searching area. If it falls on the minimum or maximum values, we create a new searching area based on the original site to ensure the selected $\alpha$ is genuinely the best.
This approach allows us to utilize feature selection without additional knowledge of the original data and modifications, as the procedure automatically finds the best alpha values, manages class balance and imbalance via under-sampling and repeated executions, and processes both perfectly balanced and highly imbalanced datasets without information loss.
Model Fine-Tuning
Having reduced the problem’s dimensionality through feature selection, we now focus on identifying the model and its configuration (hyper-parameter set) that can provide accurate predictions over unseen data samples. We selected three well-known machine learning classification models: SVM (Boser, Guyon, and Vapnik 1992), RandomForest (Breiman 2001), and state-of-the-art XGBoost (Extreme Gradient Boosting) (Chen et al. 2015).
For each model, we defined large hyper-parameter ranges to cover a variety of configurations. To reduce complexity and execution time, we developed a sampling method that randomly selects $C$ unique configurations per model (in our case, $C=50$). This method allows us to estimate different possible configurations without sacrificing execution time. Additionally, since the dataset may be affected by class imbalance, we compute class-specific weights to construct selected classifiers with consideration of class imbalance.
To evaluate each selected configuration, we use a stratified K-Fold cross-validation approach, allowing us to utilize the entire train/validation data portion and evaluate each configuration over multiple validation data portions. During the evaluation, we compute the average configuration performance measured in the F1-score, which can effectively measure model performance over both binary and multi-class classification tasks, considering class imbalance.
Afterward, we select the best classifier and configuration based on the average validation (F1-score). This model configuration becomes the final model, which will be evaluated over the hold-out (testing) data portion. In the final stage of the model development, we provide model explainability using the SHAP game-theoretical approach (Shapley 1953). This allows us to describe the reasons behind the model’s decisions and highlight the distinctions between classes.
Additionally, in the case of binary classification, we need to adjust the decision threshold, as binary classification models are not perfectly aligned with the 50% decision threshold. To do this, we utilize the precision vs. recall curve to identify the decision threshold that maximizes the model’s performance. The developed model is then trained over the entire dataset, including the train/validation and testing sets, and can be utilized for real-world applications with the identified decision threshold.
Feature Type Feature Type calculation
name_len count screen_name_sim real-valued Jaccard of (screen_name, user_name)
screen_name_len count foll_friends real-valued follower / friends
description_len count age real-valued created_at / collection date
listed count listed_by_age real-valued listed / account age
statuses count statuses_by_age real-valued statuses / account age
followers count followers_by_age real-valued followers / account age
following count following_by_age real-valued friends / account age
name_upper_len count name_upper_pcnt real-valued percentage of upper case
name_lower_len count name_lower_pcnt real-valued percentage of lower case
name_digits_len count name_digits_pcnt real-valued percentage of digits
name_special_len count name_special_pcnt real-valued percentage of other characters
screen_name_upper_len count screen_name_upper_pcnt real-valued percentage of upper case
screen_name_lower_len count screen_name_lower_pcnt real-valued percentage of lower case
screen_name_digits_len count screen_name_digits_pcnt real-valued percentage of digits
screen_name_special_len count screen_name_special_pcnt real-valued percentage of other characters
description_upper_len count description_upper_pcnt real-valued percentage of upper case
description_lower_len count description_lower_pcnt real-valued percentage of lower case
description_digits_len count description_digits_pcnt real-valued percentage of digits
description_special_len count description_special_pcnt real-valued percentage of other characters
description_urls count name_entropy real-valued entropy of user name
description_mentions count screen_name_entropy real-valued entropy of the screen name
description_hashtags count has_location boolean
total_urls count has_profile_image boolean
protected boolean has_profile_url boolean
verified boolean
Table 2: List of the profile features extracted from the Twitter API user objects.
Method Type C-15 G-17 C-17 M-18 C-S-18 C-R-19 B-F-19 TB-20 TB-22 Average
SGBot F 77.9 72.1 94.6 99.5 82.3 82.7 49.6 84.9 36.6 75.57
Kudugunta et al. F 75.3 49.8 91.7 94.5 50.9 49.2 49.6 47.3 51.7 62.22
Hayawi et al. F 85.6 34.7 93.8 91.5 60.8 60.9 20.5 77.1 24.7 61.06
BotHunter F 97.2 69.2 91.6 99.6 82.2 82.9 49.6 79.1 23.5 74.98
NameBot F 83.4 44.8 85.7 91.6 61.1 67.5 38.5 65.1 0.5 59.80
Abreu et al. F 76.4 66.7 95.0 97.9 76.9 83.5 53.8 77.1 53.4 75.63
BotArtist F 98.3 76.1 97.0 99.7 80.6 88.3 68.4 82.2 58.2 83.19
Cresci T 1.17 - 22.8 - - - - 13.7 - -
Wei T 82.7 - 78.4 - - - - 57.3 53.6 -
BGSRD T 90.8 35.7 86.3 90.5 58.2 41.1 13.0 70.0 21.1 56.30
RoBERTa T 95.8 - 94.3 - - - - 73.1 20.5 -
T5 T 89.3 - 92.3 - - - - 70.5 20.2 -
Efthimion FT 94.1 5.2 91.8 95.9 68.2 71.7 0.0 67.2 27.5 57.95
Kantepe FT 78.2 - 79.4 - - - - 62.2 58.7 -
Miller FT 83.8 59.9 86.8 91.1 56.8 43.6 0.0 74.8 45.3 60.23
Varol FT 94.7 - - - - - - 81.1 27.5 -
Kouvela FT 98.2 66.6 99.1 98.2 80.4 81.1 28.1 86.5 30.0 74.24
Santos FT 78.8 14.5 83.0 92.4 65.2 75.7 21.0 60.3 - -
Lee FT 98.6 67.8 99.3 97.9 82.5 82.7 50.3 80.0 30.4 76.61
LOBO FT 98.8 - 97.7 - - - - 80.8 38.6 -
Moghaddam FG 73.9 - - - - - - 79.9 32.1 -
Alhosseini FG 92.2 - - - - - - 72.0 38.1 -
Knauth FTG 91.2 39.1 93.4 91.3 94.0 54.2 41.3 85.2 37.1 69.64
FriendBot FTG 97.6 - 87.4 - - - - 80.0 - -
SATAR FTG 95.0 - - - - - - 86.1 - -
Botometer FTG 66.9 77.4 96.1 46.0 79.6 79.0 30.8 53.1 42.8 63.5
Rodrifuez-Ruiz FTG 87.7 - 85.7 - - - - 63.1 56.6 -
GraphHist FTG 84.5 - - - - - - 67.6 - -
EvolveBot FTG 90.1 - - - - - - 69.7 14.1 -
Dehghan FTG 88.3 - - - - - - 76.2 - -
GCN FTG 97.2 - - - - - - 80.8 54.9 -
GAT FTG 97.6 - - - - - - 85.2 55.8 -
HGT FTG 96.9 - - - - - - 88.2 39.6 -
SimpleHGN FTG 97.5 - - - - - - 88.2 45.4 -
BotRGCN FTG 97.3 - - - - - - 87.3 57.5 -
RGT FTG 97.8 - - - - - - 88.0 42.9 -
Table 3: The performance of each selected bot detection model, as reported in the (Feng et al. 2021b) paper, is compared with that of BotArtist. Performance is measured using the F1-score. In this benchmark, each model is trained and tested on each dataset separately.
Method C-15 G-17 C-17 M-18 C-S-18 C-R-19 B-F-19 TB-20 TB-22 Total Average
BotArtist 82.7 39.9 87.3 99.0 80.6 73.8 16.6 80.3 56.9 63.7 68.5
Lee 82.3 0.0 83.6 97.7 78.2 67.7 20.0 8.5 42.4 52.9 53.3
Abreu 84.4 0.3 80.1 88.4 67.1 40.8 11.7 15.6 29.0 40.4 46.3
SGBot 75.0 3.6 79.8 99.2 76.7 68.9 0.0 15.2 43.3 53.8 51.5
BotHunter 73.4 7.1 76.0 99.2 76.0 44.8 11.1 14.7 28.0 43.1 47.8
Kouvela 95.5 20.4 94.7 98.1 78.4 71.4 21.0 28.5 36.0 52.0 60.5
Botometer 66.9 77.4 96.1 46.0 79.6 79.0 30.8 53.1 42.8 45.3 63.5
Table 4: The measurement of usage, like general bot detection approaches, involves training and testing models on all datasets. Performance is assessed using the F1-score.
Feature Extraction
As mentioned earlier, in this study we pursue the creation of a simple model that in comparison with other bot detection systems does not require a large volume of information or features for accurate prediction. In our case, the number of Twitter profiles triggers our attention. The majority of social media bot accounts are created for content promotion and increasing the attention of other registered users. To maximize this goal, bot accounts pursue the expansion of social audience. Social expansion is very difficult to achieve without high activity, such as tweets, re-tweets, comments, etc. In comparison with profile-based features, the extraction and utilization of textual or graph features can result in instability, causing a significant performance drop across various periods or discussion topic datasets (Shevtsov et al. 2023a). Consequently, we need to extract as much information as possible from the user profiles to identify extremely active accounts. The difference between very popular social media accounts (also known as celebrities), and bot accounts is that they try to post and share as much content as possible during the day. To properly separate normal and bot accounts, we develop 49 unique features described in table: 2. Due to the differences between Twitter API v1.1 and v2, we have devised a method for constructing features in a way that allows the extraction of selected information from both v1.1 and v2 without any additional manipulation. The profile features we have extracted are categorized into three distinct groups: count, boolean, and real-valued.
The initial count feature categories encompass raw Twitter API object values, including followers, followings, status, and number of subscribed Twitter lists. In addition to these raw values, we have calculated statistics for textual fields such as user name, screen_name, and description. For these features, we have quantified the number of distinct characters, including uppercase letters, lowercase letters, digits, and special characters that do not fall into any of the previous categories. Furthermore, we have leveraged the profile description field, where users can provide additional information such as mentions of other users/organizations, hashtags, and additional URLs. In this context, we have counted the volume of the mentioned entities in the description field as raw values.
In contrast to the count categories, where values are represented as integers, the real-valued category contains floating-point values. We have conducted various comparisons and measurements across the profile fields in this category. Initially, we compute the age of the accounts, which allows us to distinguish highly active users with millions of posts spanning several years from users with similar activity but with an active period of only a few days. Account age is measured in days, calculated as the duration between the account creation date, provided by the Twitter API, and the collection date. Based on the account’s age, we can compute metrics similar to those in the count values category but adjusted for the account’s age perspective. For example, we compute metrics such as statuses_by_age, follower_by_age, and following_by_age using the formula:
$feature\_by\_age=\frac{feature\ value}{profile\ age}$
Given that many bot accounts are mass-created by automated scripts, they often have very similar or even trivial user names and screen names. Therefore, we have also applied the Jaccard similarity measure between user names and screen names. Jaccard similarity calculates the intersection of characters used in both fields in relation to the differences between them, using the following formula:
$Jaccard\ similarity(A,B)=\frac{|A\cap B|}{|A\cup B|}$
Additionally, we have computed the entropy of the user name and screen_name, enabling us to assess the randomness of these selected fields. In addition to the count values derived from the profile text fields (user name, screen_name, and description), we have computed the percentage of each specific character category, including lowercase, uppercase, digits, and special characters, relative to the length of the text field. These measurements offer a more precise means of identifying textual preferences more flexible to varying text lengths since they are computed as percentages of the field length.
The final category of extracted features consists of Boolean values, used to identify information that may or may not be provided by the account creator. For instance, we determine if the user account is protected, and verified, as well as extract information about the provided user location, profile URL, and whether the profile uses the default profile image.
By employing this comprehensive feature engineering approach, we have maximized feature extraction while working with the limited profile information available through Twitter API v2. In total, we have extracted 49 unique profile features, representing the full set of features used before the feature selection process in the BotArtist model.
Experimental Results
According to the methodology of SAMLP described earlier, we’ve successfully developed a semi-automated machine learning pipeline capable of constructing classification models (both binary and multi-class) for both balanced and imbalanced datasets. Using this pipeline, we created and tested our custom implementation called BotArtist for Twitter bot detection, utilizing nine well-known datasets. To evaluate BotArtist’s performance, in comparison to 34 other existing research approaches, we employed the most comprehensive Twitter bot detection benchmark available (Feng et al. 2022b), where various methods are compared using identical training/validation and testing data portions. This bench-marking allowed us to make a fair comparison between our approach and established bot detection methods.
In our comparison, we not only evaluate BotArtist against bot detection approaches based on similar categories of features but also compare it to existing approaches developed based on different sets of characteristics. The selected methods are categorized into five different groups.
Profile-Based features (F)
This category includes approaches that rely on features extracted from user profile objects, including SGBot (Yang et al. 2020), Kudugunta (Kudugunta and Ferrara 2018), Hayawi (Hayawi et al. 2022), BotHunter (Beskow and Carley 2018), NameBot (Beskow and Carley 2019) and Abreu (Abreu, Ralha, and Gondim 2020) implementations.
Text-Based features (T)
In this category, we have approaches that focus on text characteristics, including Cresci (Cresci et al. 2016), Wei (Wei and Nguyen 2019), BGSRD (Guo et al. 2021a), RoBERTa (Liu et al. 2019) and T5 (Raffel et al. 2020).
Profile and Text-Based features (FT)
These approaches combine both profile and text-based features, and the selected implementations are Efthimion (Efthimion, Payne, and Proferes 2018), Kantepe (Kantepe and Ganiz 2017), Miller (Miller et al. 2014), Varol (Varol et al. 2017), Kouvela (Kouvela, Dimitriadis, and Vakali 2020), Lee (Lee, Eoff, and Caverlee 2011) and LOBO (Echeverria et al. 2018).
Profile and Graph-Based features (FG)
Some approaches combine profile features with graph-based information. This category includes Moghaddam (Moghaddam and Abbaspour 2022) and Alhosseini (Ali Alhosseini et al. 2019) implementations.
Profile, Text, and Graph-Based features (FTG)
Approaches in this category aim to extract and combine as many features as possible, encompassing profile, text, and graph-based information. The selected implementations are Knauth (Knauth 2019), FriendBot (Beskow and Carley 2020), SATAR (Feng et al. 2021a), Botometer (Yang, Ferrara, and Menczer 2022), Rodríguez-Ruiz (Rodríguez-Ruiz et al. 2020), GraphHist (Magelinski, Beskow, and Carley 2020), EvolveBot (Yang, Harkreader, and Gu 2013), Dehghan (Dehghan et al. 2023), GCN (Kipf and Welling 2016), GAT (Veličković et al. 2017), HGT (Hu et al. 2020), SimpleHGN (Lv et al. 2021), BotRGCN (Feng et al. 2021c) and RGT (Feng et al. 2022a).
Models comparison
Our goal with BotArtist was to create a model with the highest possible level of generalizability. To achieve this, we designed two separate comparison scenarios. Initially, our focus was on assessing the fine-tuned BotArtist’s ability to capture general patterns within each dataset in our collection. As a result, we conducted training and testing for each of the selected models, including BotArtist, on each of the nine chosen datasets individually. This comparative approach allowed us to evaluate how well these selected models performed on each specific dataset, serving as a more rigorous bot detection implementation.
Additionally, the designed experiments provided us with a broader understanding of the overall capabilities of the selected methods when applied in more general scenarios. While implementing these experiments, we took into account that the performance would likely be higher than that of a generic approach. Therefore, our primary focus was on discerning the differences between the models to select the most accurate bot detection methods for the broader, more generalized comparison.
Previously, we had evaluated and compared models on each isolated dataset and calculated their average performance. While this approach provided us with an initial insight into model performance across diverse datasets, it lacked an assessment of model generalizability. To address this, we designed additional comparisons where we assessed the model’s ability to capture general patterns across different datasets. In these cases, we combined the training data from the nine datasets into a single dataset for model training. Combining data from multiple datasets, each with varying periods, discussion topics, and communities, demanded a high level of model generalization to avoid overfitting.
To measure the performance of the trained methods, we evaluated them by testing on portions of each dataset and also on the union of all the dataset testing portions. These approaches allowed us to not only identify the general performance of each bot detection method but also gain insights into performance variations across each specific dataset.
Parameter Value
max_depth
learning_rate
subsample
colsample_bytree
min_child_weight
gamma
reg_lambda
n_estimators
eval_metric
tree_method
predictor
objective
scale_pos_weight
selected features
Table 5: The following are the selected parameters of the final general BotArtist (XGBoost) model, along with the number of features chosen during the feature selection procedure.
In the initial model comparison experiment, we compared selected methods and assessed their performance based on the F1-score. This choice of metric was motivated by the highly imbalanced nature of the datasets, where proper comparison necessitates measurement metrics that account for class imbalances.
In table: 3, we present the results of this comparison. Notably, the BotArtist model emerged as the only bot detection method that achieved superior performance across three different datasets (M-18, C-R-19, B-F-19). Furthermore, the BotArtist model demonstrated the highest average F1-score, reaching 83.19. This represents a substantial improvement of 6.5% over the most accurate of the existing methods, showcasing the model’s capacity to accurately capture patterns within each distinct dataset.
As shown in table: 3, some datasets posed challenges for more complex methods that rely on text and graph features. These methods struggled to utilize datasets without text or graph information due to their limitations in accessing essential information required for the model.
In the second model comparison scenario, our objective was to assess the generalization capabilities of various models. To achieve this, we selectively choose the most accurate models from table: 3, in addition to the Botometer model (Yang, Ferrara, and Menczer 2022). The Botometer model was included because it had already been trained on a diverse range of datasets and was expected to perform well in general-case scenarios. Additionally, the SAMLP pipeline manages to fine-tune BotArtists with the parameters presented in the table: 5 where the XGBoost classification model achieves higher performance during the K-Fold cross-validation.
The results, as presented in table: 4, indicate that the most generalizable models are BotArtist, Botometer, and SGBot. BotArtist demonstrated superior performance over existing methods in terms of both F1-total and average scores, surpassing almost 10% existing methods in total. These results confirm that fine-tuned models which are based on the limited set of features are capable of capturing the general differences between normal and bot accounts.
Furthermore, it’s noteworthy that the BotArtist model outperformed existing methods in both comparison scenarios. Due to the high volume of experiments, we do not provide any information according to the selected features and final model configuration for each evaluation scenario. These data will be shared and described in detail, in a GitHub repository after the paper acceptance.
Figure 2: BotArtist model decision explainability figure produced by SHAP methodology.
Explainability
Besides the creation of the high-performance model used for bot detection, we aim to shed light on the model decision itself. For this purpose we employ SHapley Additive exPlanations (SHAP) values (Lundberg and Lee 2017) which offer various benefits. One of the most significant advantages of SHAP values is that they are not tied to any particular type of machine learning model and are therefore model-agnostic. Furthermore, SHAP values display local accuracy, and consistency, which are characteristics that are not found concurrently in other methods.
In this section we explain the BotArtist model trained over multiple datasets. Figure: 2 presents the 20 most important features and their impact on the final decision. The features are sorted in descending order of importance. Based on this order, we observe that the age of the account and the level of social activity are the most critical characteristics of our model, since the creation date of the account.
The presented figure showcases the values of the features, ranging from low (indicated by the blue color), to high (indicated by the red color). Simultaneously, the x-axis provides information on the impact of each feature value on the model decision. Based on the presented results, it is crucial to highlight that the SHAP explanation technique provides insights that align with past related works. For instance, we find that bot accounts tend to have a significantly lower age of the account and lower number of followers combined with a higher number of followers by age of the account. Based on this information we can assume that bot accounts are not so popular in terms of total followers in comparison with normal users, but the growth of the profile followers is higher in comparison with normal users. This assumption agrees with the original target of the bot account creators to achieve a higher audience during the short period.
Conclusions and Future Work
In this research paper, we introduce a semi-automatic machine learning pipeline (SAMLP) that addresses multiple challenges in creating machine learning models, including feature selection, hyperparameter fine-tuning, model evaluation, decision threshold optimization for binary classification, and provides model explainability through the SHAP game-theoretical approach. We apply this developed pipeline to create BotArtist, a versatile bot detection model based on user profile features. Our approach is trained and evaluated alongside current state-of-the-art solutions using nine different datasets. As demonstrated in our experiments, BotArtist surpasses existing, more complex methods in terms of generalization, achieving almost a 10% increase in total F1-score in multiple dataset comparisons and a 6.5% increase in separated dataset comparisons. Additionally, we offer insights into the final model’s decision-making process and the patterns it captures, based on the SHAP game-theoretical approach. Upon acceptance of this paper, we will make the developed SAMLP pipeline available for further research, along with the BotArtist model, which can label existing profile data. One of the limitations of the presented methodology is the relatively limited feature set, which might make it susceptible to evasion by future bot accounts that specifically aim to avoid detection based on these features. Further research conducted over a longer period is required to assess the model’s effectiveness in detecting bot accounts in the evolving landscape of social media.